# HuRef: HUman-REadable Fingerprint for Large Language Models

### v1

@@ 6 Conclusion
In this paper, we introduce a novel approach that uses model parameters to yield a fingerprint for LLMs. This approach could provide a human-readable identity for LLM without exposing its parameters or interfering with its training process. Different from the well-studied black-box or white-box settings, our method lies in the middle of both. Although generating the invariant terms needs access to the model parameters, which has to be done by the LLM owners and might need regulations to prevent fake terms, all the subsequent steps could be done by the public, without requiring a special third-party institution. Moreover, as only the invariant terms and fingerprint images are published, no model weights need to be exposed outside of its owner, neither to the public nor the other LLM owners, during the whole process.
Appendices
Appendix A Implementation Details of the Fingerprinting Model
A.1 Training Settings
@@
In the training stage, we alternate training the discriminator and CNN encoder every 10 steps. We set the batch size to 10, the initial learning rate to 0.0001, and introduce a noise intensity $\alpha$ of 0.16 for positive samples. After 16 epochs of training, we obtained the CNN encoder used in our paper.
A.2 Model Architecture
@@
For the CNN encoder: The CNN encoder takes invariant terms ${\bm{M}}\in\mathbb{R}^{4096\times 4096\times 6}$ as input and produces a feature vector ${\bm{v}}$ as output. Our CNN encoder structure, as depicted in Figure 4, consists of the first four convolutional layers and the last mean pooling layer. The mean pooling layer simply calculates the average of the feature maps obtained from each channel, resulting in a feature vector ${\bm{v}}$ with a length equal to the number of channels. The hyperparameters for the four convolutional layers are provided in the table below:

### v3

@@ 6 Conclusion
In this paper, we introduce a novel approach that generates a human-readable fingerprint for LLM. Owing to Zero-Knowledge Proof, all fingerprinting steps are internally conducted by the LLMs owners. Our method is actually a black-box method as only the image fingerprint and corresponding proof need to be released. There is no exposure of model weights or information leakage to the public throughout the entire process. Furthermore, we detailed our works’ limitations in Appendix K.
7 Acknowledgements
We would like to thank Shiyu Liang, Siyuan Huang, and the anonymous reviewers for helpful discussions and feedback. This work was sponsored by the National Key Research and Development Program of China (No. 2023ZD0121402) and National Natural Science Foundation of China (NSFC) grant (No.62106143).
Appendix A Additional related works
Deep neural network copyright protection methods can generally be divided into two categories: Watermarking and Fingerprinting.
Watermarking. There are three main types of watermarking methods. The first type involves embedding watermarks into model weights (Chen et al., 2019a; Wang & Kerschbaum, 2021; Liu et al., 2021; Uchida et al., 2017), hidden-layer activations (Rouhani et al., 2019), gradients (Li et al., 2022b, a), model structures (Lou et al., 2022; Chen et al., 2021), or extra components (Fan et al., 2019, 2021). These methods are typically white-box approaches and may potentially degrade the model’s performance. The second type achieves watermarking by injecting triggers into the model to produce predefined outputs (Adi et al., 2018b; Guo & Potkonjak, 2018; Le Merrer et al., 2020; Chen et al., 2019b); however, these methods often require fine-tuning or retraining the model. The third type relies on extractor subnetworks (Wu et al., 2020; Abdelnabi & Fritz, 2021) or predefined rules (He et al., 2022a, b) to embed watermarks into the model’s output.
Fingerprinting. Fingerprinting methods are primarily divided into two categories. The first category involves copyright verification through the comparison of model weights (Jia et al., 2021) or their corresponding hash values (Zheng et al., 2022; Chen et al., 2022; Xiong et al., 2022), but these methods are limited to white-box scenarios and have only been tested on CNN-based visual models. The second category includes more recent works (Zhao et al., 2020; Pan et al., 2022; Yang et al., 2022; Lukas et al., 2019; Peng et al., 2022) that construct DNN fingerprints by analyzing model behaviors on preset test cases. However, these methods often require additional models or data samples and may involve extra training.
Appendix B Details of Data Synthesis and Encoder Training
B.1 Data Synthesis
B.2 Training the Encoder
Note that we don’t need to use any real LLM weights for training the encoder, as it only needs to learn a locality-preserving mapping between the input tensor and the output Gaussian vector. This ensures strict exclusivity between the training and test data. To construct the training data, we synthesize the matrix in each channel of ${\bm{M}}$ on-the-fly, by randomly sampling 3 matrices ${\bm{P}}_{1},\displaystyle{\bm{P}}_{2},\displaystyle{\bm{P}}_{3}$ and multiplying them together as ${\bm{P}}_{1}\displaystyle{\bm{P}}_{2}\displaystyle{\bm{P}}_{3}\displaystyle{\bm{P}}_{1}^{T}$, as though they are model parameters.
To learn locality-preserving mapping, we adopt contrastive learning. For a randomly sampled input ${\bm{M}}$, its negative sample is given by another independently sampled tensor ${\bm{M}}^{-}$. For its positive sample ${\bm{M}}^{+}$, we perturb the content in each of ${\bm{M}}$’s channel by adding small perturbation noises $\epsilon_{i}\in\mathcal{N}(0,\alpha)$ to the 3 matrices behind it. Here $\alpha$ is a hyperparameter determining the small variance. (c.f. Appendix B.1 for detailed data synthesis process.)
Subsequently, the contrastive loss $\mathcal{L}_{C}$ is given by:
To render the output vector to be Gaussian, we adopt the standard GAN (Karras et al., 2019) training scheme. We add a simple MLP as the discriminator $D$ that is trained to discriminate between real Gaussian vectors and the output vector ${\bm{v}}$. In this setting, the encoder serves as the generator. During training, for every $m$ step, we alternate between training the discriminator and the generator. The discriminator loss $\mathcal{L}_{D}$ is thus given by
$\mathcal{L}=\displaystyle\mathcal{L}_{C}+\mathcal{L}_{D}$ (15)
Appendix C Implementation Details
C.1 Training Settings
@@
In the training stage, we alternate training the discriminator and encoder every 10 steps. We set the batch size to 10, the initial learning rate to 0.0001, and introduce a noise intensity $\alpha$ of 0.16 for positive samples. After 8 epochs of training, we obtained the encoder used in our paper.
C.2 Model Architecture
@@
For the encoder: We used a convolution neural network (CNN) as the encoder. The CNN encoder takes invariant terms ${\bm{M}}\in\mathbb{R}^{4096\times 4096\times 6}$ as input and produces a feature vector ${\bm{v}}$ as output. Our CNN encoder structure, as depicted in Figure 4, consists of the first four convolutional layers and the last mean pooling layer. The mean pooling layer simply calculates the average of the feature maps obtained from each channel, resulting in a feature vector ${\bm{v}}$ with a length equal to the number of channels. The hyperparameters for the four convolutional layers are provided in the table below: